# Deep DeePC: Data-enabled predictive control with low or no online optimization using deep learning

### v2

@@ 7.1 Process description
    $x_{k+1}(i)$ $=x_{k}(i)+\left(-\gamma_{i}x_{k}(i)+\frac{a_{i}}{K_{i}+x_{k}^{2}(j)}+u_{k}(i)\right)\cdot\Delta+\xi_{k}(i),\ (i,j)=\{(1,6);(2,4);(3,5)\}$ (25a)
    $x_{k+1}(i)$ $=x_{k}(i)+\left(-c_{j}x_{k}(i)+\beta_{j}x_{k}(j)\right)\cdot\Delta+\xi_{k}(i),\ (i,j)=\{(4,1);(5,2);(6,3)\}$ (25b)
@@
    7.2 Settings
    7.2.1 Control methods being evaluated
@@ 7.2.2 Parameters
    7.2.3 Simulation setting
    In this study, the following system parameters are adopted for the GRN process: $K_{i}=1$, $a_{i}=1.6$, $\gamma_{i}=0.16$, $\beta_{i}=0.16$, $c_{i}=0.06$ $(i=1,2,3)$, $\delta=0$, and $\Delta=1$ min, which are the same as those in [50]. First, open-loop simulations are conducted using the first-principles model (25) to generate data. The generated data are used for two purposes: 1) to construct the Hankel matrices $U_{p},U_{f},Y_{p},$ and $Y_{f}$, and 2) to train the DeePC operator within the proposed framework. $T=200$ is used as we construct the Hankel matrices. The data size for training the neural network is $10^{4}$. The control inputs $u(i),i=1,2,3$, are generated randomly following a uniform distribution with a prescribed range. Particularly, $u(i)\in[0,1],i=1,2,3$, and each of the inputs varies every 30 sampling periods.
    7.3 Control performance
@@
    8 Case study on chemical process
@@
    8.1 Description of the process
    This reactor-separator process consists of two continuous stirred reactors (CSTRs) and one flash tank separator. A schematic diagram of this process is presented in Figure 5. This process involves two irreversible reactions: reactant $\mathbf{A}$ converts into the desired product $\mathbf{B}$, and $\mathbf{B}$ converts into the undesired side product $\mathbf{C}$ simultaneously.
    In this process, the three vessels are interconnected through mass and energy flows. In the first reactor (CSTR 1), the input is pure reactant $\mathbf{A}$ at temperature $T_{10}$ and volumetric flow rate $F_{10}$. In the second reactor (CSTR 2), the input includes the output stream from CSTR 1 and another stream containing pure $\mathbf{A}$ at temperature $T_{20}$ and flow rate $F_{20}$. The effluent of CSTR 2 is sent to the separator with temperature $T_{2}$ and flow rate $F_{2}$. The state variables of the reactor separator process include the mass fractions of reactant $\mathbf{A}$ ($x_{\mathbf{A}i},i=1,2,3$), the mass fractions of the desired product $\mathbf{B}$ ($x_{\mathbf{B}i},i=1,2,3$), and the temperature ($T_{i},i=1,2,3$) in three vessels. The control inputs are the heat input rate $Q_{i},i=1,2,3$, to the three vessels. Typically, only the temperature $T_{i},i=1,2,3$, in three vessels are measured online using sensors. A detailed process description and the first-principles model of this process, which is used as the process simulator, can be found in [51, 21]. The objective is to implement the proposed Deep DeePC approach to drive the process operation towards desired set-points by adjusting the heat input $Q_{i},i=1,2,3$, to the three vessels.
@@ 8.1 Description of the process
    8.2 Problem formulation and simulation settings
    8.2.1 Evaluated control methods
@@ 8.2.2 Parameter settings
    8.2.3 Data generation and process setting
@@ 8.3 Results
    Figure 6 presents the closed-loop output trajectories and input trajectories obtained based on the Deep DeePC-I, constrained Deep DeePC-I, and conventional DeePC. Figure 6(a) presents the closed-loop trajectories of the output measurements based on the three controllers. In addition, the output trajectories based on open-loop control, with inputs being made the same as the steady-state reference inputs, are provided for comparison. Figure 6(b) presents the input trajectories by different controllers. All three DeePC designs can track the set-points and maintain the operation level close to the set-points. The proposed Deep DeePC-I and constrained Deep DeePC-I outperform the conventional DeePC by converging faster to the set-points. As shown in Figure 6(b), the control actions provided by Deep DeePC-I (shown in red lines) sometimes violate the input constraints. With the implementation of the proposed constraint handling scheme, constrained Deep DeePC-I (shown in blue dashed-dotted lines) ensures that the system constraints are satisfied throughout the entire operation.
@@
    In real-world applications, it can be challenging or costly to measure certain state variables using hardware sensors in real-time. For example, measuring concentrations of substances often requires a lab analyzer, which can be time-consuming and presents challenges for implementing online control methods. It is worth noting that the proposed Deep DeePC method, owing to the adoption of the DeePC framework, requires only system output feedback (specifically, temperature measurements in the reactor-separator process). This approach offers enhanced practicality compared to optimal control methods that require full-state feedback.
    8.4 Comparisons of computation time
@@
    We also assess how the parameters affect the computation time of conventional DeePC. Table 10 presents the average computation times for DeePC with different values of $T$. The results indicate that the computation time for the DeePC optimization problem increases rapidly as $T$ increases. This suggests that the complexity of the conventional DeePC optimization problem in (12) is significantly influenced by the parameters $T$, $T_{ini}$, and $N_{p}$. To satisfy the persistently exciting condition described in Definition 1, $T$ generally needs to be made sufficiently large, which leads to increased computation time for DeePC. In contrast, the proposed Deep DeePC methods are minimally affected by these parameters ($T$, $T_{ini}$, and $N_{p}$), since either no optimization or only low optimization needs to be solved during online implementation.
    9 Conclusion
    In this paper, we proposed a computation-efficient, deep learning-enabled DeePC control approach for nonlinear processes, referred to as Deep DeePC. Within the proposed framework, a deep neural network is used to output the DeePC operator $g$ at each sampling instant, and the generated DeePC operator is then used to generate an optimal control input sequence. The parameters of this neural network can be trained using input and output data collected from open-loop process operations. During online implementation, the Deep DeePC framework leverages the trained neural network to directly output the DeePC operator. Therefore, online optimization required by conventional DeePC is bypassed. To further address constrained control problems, a constraint handling scheme was developed and integrated with the proposed Deep DeePC. The proposed approach was applied to both a gene regulatory network and a reactor-separator chemical process, good control performance was achieved in both case studies. The results demonstrated that this method can drive process operations toward desired set-points even when the reference inputs corresponding to those set-points are not provided. Additionally, the online computation time for executing the proposed Deep DeePC is significantly reduced compared to conventional DeePC. The constrained Deep DeePC can handle input and control constraints, while maintaining more efficient online computation as compared to conventional DeePC.
@@ Acknowledgment
    This research is supported by the National Research Foundation, Singapore, and PUB, Singapore’s National Water Agency under its RIE2025 Urban Solutions and Sustainability (USS) (Water) Centre of Excellence (CoE) Programme, awarded to Nanyang Environment & Water Research Institute (NEWRI), Nanyang Technological University, Singapore (NTU). This research is supported by Ministry of Education, Singapore, under its Academic Research Fund Tier 1 (RS15/21 & RG63/22), and Nanyang Technological University, Singapore (Start-Up Grant). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore and PUB, Singapore’s National Water Agency.
@@
    The numerical data used to generate Figures 4, 6, 7, and Tables 3, 7, 8 are provided in the Supplementary Material. The compressed file contains data obtained from the case studies on the gene regulatory network and the reactor-separator process. This encompasses simulated historical open-loop data for neural network training, maximum and minimum values of states used for scaling, and the simulated closed-loop trajectories based on different control approaches (i.e., our proposed Deep DeePC, conventional DeePC, and open-loop control). The data can be used to reproduce and generate the figures and tables presented in this paper. Specifically, the RMSEs in Tables 3, 7, and 8 can be obtained using the reference trajectories and the trajectories provided by different controllers. Additionally, the Hankel matrices and the training data for the neural network can be constructed using historical open-loop data, following the instructions provided in this paper. The source code for our methods is available at https://github.com/QiYuan-Zhang/Deep-DeePC.

### v3

@@ Introduction
    $x_{k+1}(i)$ $=x_{k}(i)+\left(-\gamma_{i}x_{k}(i)+\frac{a_{i}}{K_{i}+x_{k}^{2}(j)}+u_{k}(i)\right)\cdot\Delta+\xi_{k}(i),\ (i,j)\in\{(1,6);(2,4);(3,5)\}$ (25a)
    $x_{k+1}(i)$ $=x_{k}(i)+\left(-c_{j}x_{k}(i)+\beta_{j}x_{k}(j)\right)\cdot\Delta+\xi_{k}(i),\ (i,j)\in\{(4,1);(5,2);(6,3)\}$ (25b)
@@
    Settings
    Control methods being evaluated
@@
    Simulation setting
    In this study, the following system parameters are adopted for the GRN process: $K_{i}=1$, $a_{i}=1.6$, $\gamma_{i}=0.16$, $\beta_{i}=0.16$, $c_{i}=0.06$ $(i=1,2,3)$, $\delta=0$, and $\Delta=1$ min, which are the same as those in Reference [52]. First, open-loop simulations are conducted using the first-principles model (25) to generate data. The generated data are used for two purposes: 1) to construct the Hankel matrices $U_{p},U_{f},Y_{p},$ and $Y_{f}$, and 2) to train the DeePC operator within the proposed framework. $T=200$ is used as we construct the Hankel matrices. The data size for training the neural network is $10^{4}$. The control inputs $u(i),i=1,2,3$, are generated randomly following a uniform distribution with a prescribed range. Particularly, $u(i)\in[0,1],i=1,2,3$, and each of the inputs varies every 30 sampling periods.
@@
    Control performance
@@
    Case study on chemical process
@@
    Description of the process
    This reactor-separator process consists of two continuous stirred tank reactors (CSTRs) and one flash tank separator. A schematic diagram of this process is presented in Figure 5. This process involves two irreversible reactions: reactant $\mathbf{A}$ converts into the desired product $\mathbf{B}$, and $\mathbf{B}$ converts into the undesired side product $\mathbf{C}$, simultaneously.
    In this process, the three vessels are interconnected through mass and energy flows. In the first reactor (CSTR 1), the input is pure reactant $\mathbf{A}$ at temperature $T_{10}$ and volumetric flow rate $F_{10}$. In the second reactor (CSTR 2), the input includes the output stream from CSTR 1 and another stream containing pure $\mathbf{A}$ at temperature $T_{20}$ and flow rate $F_{20}$. The effluent of CSTR 2 is sent to the separator with temperature $T_{2}$ and flow rate $F_{2}$. The state variables of the reactor separator process include the mass fractions of reactant $\mathbf{A}$ ($x_{\mathbf{A}i},i=1,2,3$), the mass fractions of the desired product $\mathbf{B}$ ($x_{\mathbf{B}i},i=1,2,3$), and the temperature ($T_{i},i=1,2,3$) in three vessels. The control inputs are the heat input rate $Q_{i},i=1,2,3$, to the three vessels. Typically, only the temperature $T_{i},i=1,2,3$, in three vessels are measured online using sensors. A detailed process description and the first-principles model of this process, which is used as the process simulator, can be found in References [21, 53]. The objective is to implement the proposed Deep DeePC approach to drive the process operation towards desired set-points by adjusting the heat input $Q_{i},i=1,2,3$, to the three vessels.
@@
    Problem formulation and simulation settings
    Evaluated control methods
@@
    Remark 2
    In this work, we primarily focus on integrating neural networks with conventional DeePC to achieve efficient online computation. In the two case studies, we employed fully-connected feedforward deep neural networks. Each of the neural networks in the two case studies has two hidden layers. When dealing with systems with larger scales, the number of hidden layers may need to be increased.
    Remark 3
    The control performance of the proposed Deep DeePC approach is dependent on several tuning parameters, including the length of historical data, the length of initial trajectories, the length of future prediction trajectories, and weighting matrices in (14b). These parameters need to be tuned through trial-and-error. If the control performance of Deep DeePC remains unsatisfactory, it is likely due to insufficient training data for the considered system. In such cases, acquiring additional data for neural network training may be necessary.
    Data generation and process setting
@@ Results
    Figure 6 presents the closed-loop output trajectories and input trajectories obtained based on the Deep DeePC-I, constrained Deep DeePC-I, and conventional DeePC. Figure 6(a) presents the closed-loop trajectories of the controlled output based on the three controllers. In addition, the output trajectories based on open-loop control, with inputs being made the same as the steady-state reference inputs, are provided for comparison. Figure 6(b) presents the input trajectories by different controllers. All three DeePC designs can track the set-points and maintain the operation level close to the set-points. The proposed Deep DeePC-I and constrained Deep DeePC-I outperform the conventional DeePC by converging faster to the set-points. As shown in Figure 6(b), the control actions provided by Deep DeePC-I (shown in red lines) sometimes violate the input constraints. With the implementation of the proposed constraint handling scheme, constrained Deep DeePC-I (shown in blue dashed-dotted lines) ensures that the system constraints are satisfied throughout the entire operation.
@@
    Representative data-driven model predictive control methods typically utilize data information to develop a dynamic model of the system, and then formulate an online MPC optimization problem based on the data-driven model [54, 55, 56, 20, 57]. It is worth mentioning that these data-driven MPC methods typically require full-state information. In certain applications, measuring specific state variables in real-time with hardware sensors can be challenging or expensive. The Deep DeePC control scheme requires only control inputs and controlled outputs, eliminating the need to establish an explicit dynamic model. In such cases, the proposed approach can be more advantageous.
    Comparisons of computation time
@@
    We also assess how the parameters affect the computation time of conventional DeePC. Table 10 presents the average computation times for DeePC with different values of $T$. The results indicate that the computation time for the DeePC optimization problem increases rapidly as $T$ increases. This suggests that the complexity of the conventional DeePC optimization problem in (12) is significantly influenced by the parameters $T$. To satisfy the persistently exciting condition described in Definition 1, $T$ generally needs to be made sufficiently large, which leads to increased computation time for DeePC. In contrast, the proposed Deep DeePC methods are minimally affected by these parameters ($T$, $T_{ini}$, and $N_{p}$), since either no optimization or only low optimization needs to be solved during online implementation.
    Conclusion
    In this paper, we proposed a computation-efficient, deep learning-enabled DeePC control approach for nonlinear processes, referred to as Deep DeePC. Within the proposed framework, a deep neural network was used to output the DeePC operator $g$ at each sampling instant, and the generated DeePC operator was then used to generate an optimal control input sequence. The parameters of this neural network can be trained using input and output data collected from open-loop process operations. During online implementation, the Deep DeePC framework leverages the trained neural network to directly output the DeePC operator. Therefore, online optimization required by conventional DeePC is bypassed. To further address constrained control problems, a constraint handling scheme was developed and integrated with the proposed Deep DeePC. The proposed approach was applied to both a gene regulatory network and a reactor-separator chemical process, good control performance was achieved in both case studies. The results demonstrated that this method can drive process operations toward desired set-points even when the reference inputs corresponding to those set-points are not provided. Additionally, the online computation time for executing the proposed Deep DeePC is significantly reduced compared to conventional DeePC. The constrained Deep DeePC can handle input and control constraints, while maintaining more efficient online computation as compared to conventional DeePC.
@@ Acknowledgment
    This research is supported by the National Research Foundation, Singapore, and PUB, Singapore’s National Water Agency under its RIE2025 Urban Solutions and Sustainability (USS) (Water) Centre of Excellence (CoE) Programme, awarded to Nanyang Environment & Water Research Institute (NEWRI), Nanyang Technological University, Singapore (NTU). This research is also supported by Ministry of Education, Singapore, under its Academic Research Fund Tier 1 (RS15/21 & RG63/22), and Nanyang Technological University, Singapore (Start-Up Grant). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore and PUB, Singapore’s National Water Agency.
@@
    The numerical data used to generate Figures 4, 6, 7, and Tables 3, 7, 8 are provided in the Supplementary Material. The compressed file contains data obtained from the case studies on the gene regulatory network and the reactor-separator process. This encompasses simulated historical open-loop data for neural network training, maximum and minimum values of states used for scaling, and the simulated closed-loop trajectories based on different control approaches (i.e., our proposed Deep DeePC, conventional DeePC, and open-loop control). The data can be used to reproduce and generate the figures and tables presented in this paper. Specifically, the RMSEs in Tables 3, 7, and 8 can be obtained using the reference trajectories and the trajectories provided by different controllers. Additionally, the Hankel matrices and the training data for the neural network can be constructed using historical open-loop data, following the instructions provided in this paper. The source code for our methods is available at https://github.com/Zhang-Xuewen/Deep-DeePC.